hi
im dawn marie
my question is
jen was 25 years old when her son was born, her present age is 7 years less than 3 times her son's present age. how old is jen now?
please answer my question...

Answer by bucky(2189)

When Jen's son was born she was 25 yrs old and her son was 0 (not yet 1). Let x be the number of
years that have gone by since then. This means that Jen is now 25 + x years old and her
son is x years old.
.
If you take 3 times her sons present age and subtract 7 from that, the result is Jen's present age.
.
We can write an equation that says exactly that: three times her son's age less 7 is 3x - 7
and this is equal to Jen's present age of 25 + x. In equation form this is:
.
3x - 7 = 25 + x
.
Get rid of the x on the right side by subtracting x from both sides and you have:
.
2x - 7 = 25
.
Get rid of the -7 on the left side by adding 7 to both sides to reduce the equation to:
.
2x = 32
.
Solve for x by dividing both sides of this equation by 2 and you have:
.
x = 16
.
16 years have gone by since her son was born. Since Jen was 25 when he was born, that means
that Jen is now 41 years old (25 + 16).
.
Check: is 41 equal to three times her son's age less 7 years. Her son is 16 and 3 times 16
is 48 and take away 7 results in 41. Works OK, so our answer is correct.
.
Hope this helps you to understand the problem.
.